a person invests $310 in an account that pays 8%. how do I find the balance after 10 years?

Answer by Theo(13342)

if you assume compound interest where the interest is compounded annually, then the investment after 10 years will be 310 * (1.08)^10 = 669.27.

if you assume simple interest, then the investment after 10 years will be 310 + 310 * .08 * 10 = 558.

if you assume compound interest where the interest is compounded monthly, then the investment after 10 years will be 310 * (1 + .08/12)^10*12 = 310 * (1.006666666...)^120 = 688.09

if you assume continuous compounding of interest, then the investment after 10 years will be 310 * e^(.08*10) = 689.92.

in other words, you need to understand how the interest is handled before determing what the value will be after 10 years.